### v1

@@ III.3 $D^{*}\bar{B}^{*}$ systems
$\mathcal{M}_{a3}^{(0)}=$ $~{}8[(D_{a}-D_{b}+E_{a}-E_{b})(\mathcal{O}_{1}+\mathcal{O}_{2})$
$+2(D_{b}+E_{b})\mathcal{O}_{3}],$ (50)
$\mathcal{M}_{b3}^{(0)}=$ $-\frac{g^{2}}{f^{2}}\frac{\mathcal{G}(q,\epsilon_{1},\epsilon_{2},\epsilon_{3}^{*},\epsilon_{4}^{*})}{q^{2}-m_{\pi}^{2}},$ (51)
@@
$\mathcal{M}_{(a3)}^{(0)}=$ $~{}8[(D_{a}-D_{b}+E_{a}-E_{b})(\mathcal{O}_{1}+\mathcal{O}_{2})$
$+2(D_{b}+E_{b})\mathcal{O}_{3}],$ (52)
$\mathcal{M}_{(b3)}^{(0)}$ $=-\frac{g^{2}}{f^{2}}\frac{\mathcal{G}(q,\epsilon_{1},\epsilon_{2},\epsilon_{3}^{*},\epsilon_{4}^{*})}{q^{2}-m_{\pi}^{2}},$ (53)
@@
$\mathcal{M}_{a_{3.1}}^{(2)}=$ $~{}-4\frac{g_{2}^{2}}{f^{2}}[(d-3)(d-2)A_{1}\mathcal{O}_{1}-(d-3)A_{2}\mathcal{O}_{2}$
@@
$\mathcal{M}_{a_{3.9}}^{(2)}=$ $4\frac{g_{1}g_{2}}{f^{2}}(d-3)[A_{1}\mathcal{O}_{1}-A_{2}\mathcal{O}_{2}+A_{3}\mathcal{O}_{3}]J^{h}_{22},$ (65)
$\mathcal{M}_{a{3.10}}^{(2)}=$ $4\frac{g_{1}^{2}}{f^{2}}A_{1}\mathcal{O}_{3}J^{h}_{22},$ (66)
$\mathcal{M}_{a{3.11}}^{(2)}=$ $4\frac{g_{1}^{2}}{f^{2}}(d-3)[A_{1}\mathcal{O}_{1}-A_{2}\mathcal{O}_{2}]J^{h}_{22},$ (67)
$\mathcal{M}_{a{3.12}}^{(2)}=$ $4\frac{g_{1}^{2}}{f^{2}}(d-3)[A_{1}\mathcal{O}_{1}-A_{2}\mathcal{O}_{2}]J^{h}_{22},$ (68)
$\mathcal{M}_{a{3.13}}^{(2)}=$ $4\frac{g_{1}g_{2}}{f^{2}}(d-3)[A_{1}\mathcal{O}_{1}-A_{2}\mathcal{O}_{2}+A_{3}\mathcal{O}_{3}]J^{h}_{22},$ (69)
$\mathcal{M}_{a_{3.14}}^{(2)}=$ $4\frac{g_{1}g_{2}}{f^{2}}A_{1}\mathcal{O}_{3}J^{h}_{22},$ (70)
$\mathcal{M}_{a_{3.15}}^{(2)}=$ $4\frac{g_{1}g_{2}}{f^{2}}(d-3)[A_{1}\mathcal{O}_{1}-A_{2}\mathcal{O}_{2}]J^{h}_{22},$ (71)
$\mathcal{M}_{a_{3.16}}^{(2)}=$ $4\frac{g_{1}g_{2}}{f^{2}}(d-3)[A_{1}\mathcal{O}_{1}-A_{2}\mathcal{O}_{2}]J^{h}_{22},$ (72)
$\mathcal{M}_{a_{3.(17+18)}}^{(2)}=$ $-\frac{3}{8}\frac{g_{2}^{2}}{f^{2}}A_{1}[(d-2)\partial_{\omega}J^{b}_{22}(\omega_{1})+\partial_{\omega}J^{b}_{22}(\omega_{2})],$ (73)
$\mathcal{M}_{a_{3.(19+20)}}^{(2)}=$ $-\frac{3}{8}\frac{g_{1}^{2}}{f^{2}}A_{1}[(d-2)\partial_{\omega}J^{b}_{22}(\omega_{1})+\partial_{\omega}J^{b}_{22}(\omega_{2})],$ (74)
@@
$\mathcal{M}_{b{3.1}}^{(2)}=$ $4\frac{g_{1}^{3}g_{2}}{f^{2}}A_{1}J^{g}_{22}\frac{\mathcal{G}(q,\epsilon_{1},\epsilon_{2},\epsilon_{3}^{*},\epsilon_{4}^{*})}{q^{2}-m_{\pi}^{2}},$ (75)
@@
$\mathcal{M}_{b{3.6}}^{(2)}=$ $4\frac{g_{1}g_{2}^{3}}{f^{2}}A_{1}J^{g}_{22}\frac{\mathcal{G}(q,\epsilon_{1},\epsilon_{2},\epsilon_{3}^{*},\epsilon_{4}^{*})}{q^{2}-m_{\pi}^{2}},$ (80)
$\mathcal{M}_{b{3.7}}^{(2)}=$ $-\frac{8}{3}\frac{g_{1}g_{2}}{f^{2}}A_{1}J^{c}_{0}\frac{\mathcal{G}(q,\epsilon_{1},\epsilon_{2},\epsilon_{3}^{*},\epsilon_{4}^{*})}{q^{2}-m_{\pi}^{2}},$ (81)
$\mathcal{M}_{b{3.8}}^{(2)}=$ $-\frac{8}{3}\frac{g_{1}g_{2}}{f^{2}}A_{1}J^{c}_{0}\frac{\mathcal{G}(q,\epsilon_{1},\epsilon_{2},\epsilon_{3}^{*},\epsilon_{4}^{*})}{q^{2}-m_{\pi}^{2}},$ (82)
$\mathcal{M}_{b{3.9}}^{(2)}=$ $-\frac{8}{3}\frac{g_{1}g_{2}}{f^{2}}A_{1}[2m^{2}L+\frac{m^{2}}{16\pi^{2}}\rm log(\frac{m^{2}}{\mu^{2}})]$
@@
$\mathcal{M}_{b{3.14}}^{(2)}=$ $\mathcal{M}_{b_{3.15}}^{(2)}=\mathcal{M}_{b_{3.16}}^{(2)}=\mathcal{M}_{b_{3.17}}^{(2)}=0.$ (86)
@@
$A_{a3.1}$ $\frac{1}{4}(3D_{a}-E_{a})$ $\frac{1}{4}(3D_{b}-E_{b})$ $\frac{1}{4}(3D_{b}-E_{b}$) $\frac{3}{4}(D_{b}+E_{b})$ $\frac{3}{4}(D_{b}+E_{b})$ $\frac{3}{4}(D_{b}+E_{b})$ 00 00
@@
$A_{a3.3(4)}$ $\frac{1}{4}(3D_{a}-E_{a})$ $\frac{1}{4}(3D_{a}-E_{a})$ 00 $\frac{3}{4}(D_{b}+E_{b})$ $\frac{3}{4}(D_{b}+E_{b})$ 00 $0(\delta_{2})$ $\delta_{2}(0)$
@@
$A_{a3.9}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{1}{4}(D_{a}+E_{a})$ $\frac{1}{4}[(d-3)(D_{b}+E_{b})$ $+(D_{a}+E_{a})]$ $\frac{-3}{4}(D_{b}-3E_{b})$ $\frac{-3}{4}(D_{b}-3E_{b})$ $\frac{-3}{4}[(d-3)(D_{b}-3E_{b})$ $+(D_{a}-3E_{a})]$ 00 00
@@
$A_{a3.13}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{1}{4}(D_{a}+E_{a})$ $\frac{1}{4}[(d-3)(D_{b}+E_{b})$ $+(D_{a}+E_{a})]$ $\frac{-3}{4}(D_{b}-3E_{b})$ $\frac{-3}{4}(D_{b}-3E_{b})$ $\frac{-3}{4}[(d-3)(D_{b}-3E_{b})$ $+(D_{a}-3E_{a})]$ 00 00
@@
       $A_{c3.13}$        $\frac{5}{16}$        00        00        $\frac{-3}{16}$        00        00        00        $\delta_{2}$
@@ IV.1 Potentials in coordinate space and possible bound states
After obtaining the scattering amplitudes, we now evaluate the effective potentials and analyze their behaviors in coordinate space. To obtain the numerical results, we use the following LECs, determined by the resonance saturation model Xu:2017tsr ; Wang:2018atz ; Xu:2021vsi :
@@
Other parameters include: $m_{\pi}=0.139~{}\rm{GeV}$, the pion decay constant $f_{\pi}=0.086~{}\rm{GeV}$, the renormalization scale $\mu=4\pi f_{\pi}$, $\delta_{1}=0.142~{}\rm{GeV}$, $\delta_{2}=0.045~{}\rm{GeV}$, and the coupling constants $g_{1}=0.065$ and $g_{2}=0.52$ Wang:2018atz .
@@ IV.1.3 $D^{*}\bar{B}^{*}$ systems
Besides, in the $D^{*}\bar{B}^{*}[0(0^{+})]$ and $D^{*}\bar{B}^{*}[0(1^{+})]$ channels, two shallow bound states are found with binding energies of $6.0~{}\rm{MeV}$ and $0.6~{}\rm{MeV}$, respectively. Therefore, we suggest that future theoretical and experimental studies should focus on these two systems to find new heavy tetraquark states. Compared the total potentials of the five $I=0$ channels, we can also conclude that the interaction in $D^{*}\bar{B}^{*}[0(0^{+})]$ channel is more attractive than the other four channels. The mass and the binding energy of each $D^{(*)}\bar{B}^{(*)}[I=0]$ system under different cutoffs are listed in Table 9.
@@
IV.2 Two-body scattering of $D^{(*)}\bar{B}^{(*)}$ systems
In addition to the binding energies, we further derive information about $D^{(*)}\bar{B}^{(*)}$ elastic scattering by solving the Lippmann-Schwinger equation. In quantum theory, the scattering rate represents the probability per unit time for a scattering event, it is proportional to the cross section, the number of the targets, and the flux Thomson:2013zua . Using Eqs. (12)-(16), we now calculate the scattering rates between $D^{(*)}$ and $\bar{B}^{(*)}$. The variations of the scattering rates $k\sigma(k)$ with cutoffs are shown in Figs. 9 and 10, as functions of the center-of-mass energy.

### v2

@@ III.3 $D^{*}\bar{B}^{*}$ systems
$\mathcal{M}_{a3}^{(0)}=$ $~{}4[(D_{a}+E_{a})\mathcal{O}_{1}-(D_{b}+E_{b})\mathcal{O}_{2}$
$+(D_{b}+E_{b})\mathcal{O}_{3}],$ (50)
$\mathcal{M}_{b3}^{(0)}=$ $-\frac{g_{1}g_{2}}{f^{2}}\frac{\mathcal{G}(q,\epsilon_{1},\epsilon_{2},\epsilon_{3}^{*},\epsilon_{4}^{*})}{q^{2}-m_{\pi}^{2}},$ (51)
@@
$\mathcal{M}_{(a3)}^{(0)}=$ $~{}4[(D_{a}+E_{a})\mathcal{O}_{1}-(D_{b}+E_{b})\mathcal{O}_{2}$
$+(D_{b}+E_{b})\mathcal{O}_{3}],$ (52)
$\mathcal{M}_{(b3)}^{(0)}=$ $-\frac{g_{1}g_{2}}{f^{2}}\frac{\mathcal{G}(q,\epsilon_{1},\epsilon_{2},\epsilon_{3}^{*},\epsilon_{4}^{*})}{q^{2}-m_{\pi}^{2}},$ (53)
@@
$\mathcal{M}_{a_{3.1}}^{(2)}=$ $-4\frac{g_{2}^{2}}{f^{2}}[(d-3)(d-2)A_{1}\mathcal{O}_{1}-(d-3)A_{2}\mathcal{O}_{2}$
@@
$\mathcal{M}_{a_{3.9}}^{(2)}=$ $~{}4\frac{g_{1}g_{2}}{f^{2}}(d-3)[A_{1}\mathcal{O}_{1}-A_{2}\mathcal{O}_{2}+A_{3}\mathcal{O}_{3}]J^{h}_{22},$ (65)
$\mathcal{M}_{a{3.10}}^{(2)}=$ $~{}4\frac{g_{1}^{2}}{f^{2}}A_{1}\mathcal{O}_{3}J^{h}_{22},$ (66)
$\mathcal{M}_{a{3.11}}^{(2)}=$ $~{}4\frac{g_{1}^{2}}{f^{2}}(d-3)[A_{1}\mathcal{O}_{1}-A_{2}\mathcal{O}_{2}]J^{h}_{22},$ (67)
$\mathcal{M}_{a{3.12}}^{(2)}=$ $~{}4\frac{g_{1}^{2}}{f^{2}}(d-3)[A_{1}\mathcal{O}_{1}-A_{2}\mathcal{O}_{2}]J^{h}_{22},$ (68)
$\mathcal{M}_{a{3.13}}^{(2)}=$ $~{}4\frac{g_{1}g_{2}}{f^{2}}(d-3)[A_{1}\mathcal{O}_{1}-A_{2}\mathcal{O}_{2}+A_{3}\mathcal{O}_{3}]J^{h}_{22},$ (69)
$\mathcal{M}_{a_{3.14}}^{(2)}=$ $~{}4\frac{g_{1}g_{2}}{f^{2}}A_{1}\mathcal{O}_{3}J^{h}_{22},$ (70)
$\mathcal{M}_{a_{3.15}}^{(2)}=$ $~{}4\frac{g_{1}g_{2}}{f^{2}}(d-3)[A_{1}\mathcal{O}_{1}-A_{2}\mathcal{O}_{2}]J^{h}_{22},$ (71)
$\mathcal{M}_{a_{3.16}}^{(2)}=$ $~{}4\frac{g_{1}g_{2}}{f^{2}}(d-3)[A_{1}\mathcal{O}_{1}-A_{2}\mathcal{O}_{2}]J^{h}_{22},$ (72)
$\mathcal{M}_{a_{3.(17+18)}}^{(2)}=$ $-\frac{3g_{2}^{2}}{2f^{2}}A_{1}[(d-2)\partial_{\omega}J^{b}_{22}(\omega_{1})+\partial_{\omega}J^{b}_{22}(\omega_{2})],$ (73)
$\mathcal{M}_{a_{3.(19+20)}}^{(2)}=$ $-\frac{3g_{1}^{2}}{2f^{2}}A_{1}[(d-2)\partial_{\omega}J^{b}_{22}(\omega_{1})+\partial_{\omega}J^{b}_{22}(\omega_{2})],$ (74)
@@
$\mathcal{M}_{b{3.1}}^{(2)}=$ $~{}4\frac{g_{1}^{3}g_{2}}{f^{2}}A_{1}J^{g}_{22}\frac{\mathcal{G}(q,\epsilon_{1},\epsilon_{2},\epsilon_{3}^{*},\epsilon_{4}^{*})}{q^{2}-m_{\pi}^{2}},$ (75)
@@
$\mathcal{M}_{b{3.6}}^{(2)}=$ $~{}4\frac{g_{1}g_{2}^{3}}{f^{2}}A_{1}J^{g}_{22}\frac{\mathcal{G}(q,\epsilon_{1},\epsilon_{2},\epsilon_{3}^{*},\epsilon_{4}^{*})}{q^{2}-m_{\pi}^{2}},$ (80)
$\mathcal{M}_{b{3.7}}^{(2)}=$ $-\frac{8g_{1}g_{2}}{3f^{2}}A_{1}J^{c}_{0}\frac{\mathcal{G}(q,\epsilon_{1},\epsilon_{2},\epsilon_{3}^{*},\epsilon_{4}^{*})}{q^{2}-m_{\pi}^{2}},$ (81)
$\mathcal{M}_{b{3.8}}^{(2)}=$ $-\frac{8g_{1}g_{2}}{3f^{2}}A_{1}J^{c}_{0}\frac{\mathcal{G}(q,\epsilon_{1},\epsilon_{2},\epsilon_{3}^{*},\epsilon_{4}^{*})}{q^{2}-m_{\pi}^{2}},$ (82)
$\mathcal{M}_{b{3.9}}^{(2)}=$ $-\frac{8g_{1}g_{2}}{3f^{2}}A_{1}[2m^{2}L+\frac{m^{2}}{16\pi^{2}}{\rm{log}}(\frac{m^{2}}{\mu^{2}})]$
@@
$\mathcal{M}_{b3.14}^{(2)}=$ $~{}\mathcal{M}_{b3.15}^{(2)}=\mathcal{M}_{b3.16}^{(2)}=\mathcal{M}_{b3.17}^{(2)}=0.$ (86)
@@
$A_{a3.1}$ $\frac{1}{4}(3D_{a}-E_{a})$ $\frac{1}{4}(3D_{b}-E_{b})$ $\frac{1}{4}(3D_{b}-E_{b}$) $\frac{3}{4}(D_{a}+E_{a})$ $\frac{3}{4}(D_{b}+E_{b})$ $\frac{3}{4}(D_{b}+E_{b})$ 00 00
@@
$A_{a3.3(4)}$ $\frac{1}{4}(3D_{b}-E_{b})$ $\frac{1}{4}(3D_{b}-E_{b})$ 00 $\frac{3}{4}(D_{b}+E_{b})$ $\frac{3}{4}(D_{b}+E_{b})$ 00 $0(\delta_{2})$ $\delta_{2}(0)$
@@
$A_{a3.9}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{1}{4}(D_{a}+E_{a})$ $\frac{1}{4}[(d-3)(D_{b}+E_{b})$ $+(D_{a}+E_{a})]$ $\frac{-3}{4}(D_{b}-3E_{b})$ $\frac{-3}{4}(D_{a}-3E_{a})$ $\frac{-3}{4}[(d-3)(D_{b}-3E_{b})$ $+(D_{a}-3E_{a})]$ 00 00
@@
$A_{a3.13}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{1}{4}(D_{a}+E_{a})$ $\frac{1}{4}[(d-3)(D_{b}+E_{b})$ $+(D_{a}+E_{a})]$ $\frac{-3}{4}(D_{b}-3E_{b})$ $\frac{-3}{4}(D_{a}-3E_{a})$ $\frac{-3}{4}[(d-3)(D_{b}-3E_{b})$ $+(D_{a}-3E_{a})]$ 00 00
@@
       $A_{c3.13}$        $\frac{5}{16}$        00        00        $-\frac{3}{16}$        00        00        00        $\delta_{2}$
@@ IV.1 Potentials in coordinate space and possible bound states
After obtaining the scattering amplitudes, we now evaluate the effective potentials and analyze their behavior in coordinate space. To obtain the numerical results, we use the following LECs, determined by the resonance saturation model Xu:2017tsr ; Wang:2018atz ; Xu:2021vsi ; Du:2016tgp ; Ecker:1988te ; Ecker:1989yg ; Donoghue:1988ed :
@@
Other parameters include: $m_{\pi}=0.139~{}\rm{GeV}$, the pion decay constant $f_{\pi}=0.086~{}\rm{GeV}$, the renormalization scale $\mu=4\pi f_{\pi}$, $\delta_{1}=0.142~{}\rm{GeV}$, $\delta_{2}=0.045~{}\rm{GeV}$, and the coupling constants $g_{1}=0.65$ and $g_{2}=0.52$ Xu:2017tsr ; Wang:2018atz .
@@ IV.1.3 $D^{*}\bar{B}^{*}$ systems
Besides, in the $D^{*}\bar{B}^{*}[0(0^{+})]$ and $D^{*}\bar{B}^{*}[0(1^{+})]$ channels, two shallow bound states are found with binding energies of $6.0~{}\rm{MeV}$ and $0.6~{}\rm{MeV}$, respectively. Therefore, we suggest that future theoretical and experimental studies should focus on these two systems to find new heavy tetraquark states. Compared to the total potentials of the five $I=0$ channels, we can also conclude that the interaction in $D^{*}\bar{B}^{*}[0(0^{+})]$ channel is more attractive than the other four channels. The mass and the binding energy of each $D^{(*)}\bar{B}^{(*)}[I=0]$ system under different cutoffs are listed in Table 9.
@@
IV.2 Coupled-channel effects
We have considered interactions of the $D^{(*)}\bar{B}^{(*)}$ systems in the single-channel case. In this section, we will investigate how much their coupled-channel effects affect the above results. We categorize $I=0$ coupled-channel $D^{(*)}\bar{B}^{(*)}$ systems into two groups according to their $0(J^{P})$ quantum number: $D\bar{B}(^{1}S_{0})$, $D^{*}\bar{B}^{*}(^{1}S_{0})$}$[0(0^{+})]$ and $D\bar{B}^{*}(^{3}S_{1})$, $D^{*}\bar{B}(^{3}S_{1})$, $D^{*}\bar{B}^{*}(^{3}S_{1})$}$[0(1^{+})]$.
We first consider the $D^{(*)}\bar{B}^{(*)}[0(0^{+})]$ system composed of $D\bar{B}(^{1}S_{0})$, $D^{*}\bar{B}^{*}(^{1}S_{0})$}. In this case, the effective potentials of the inelastic scattering process $D\bar{B}\to D^{*}\bar{B}^{*}$ and $D^{*}\bar{B}^{*}\to D\bar{B}$ are needed. We depict the $D\bar{B}\to D^{*}\bar{B}^{*}$ diagrams in Fig. 12. At LO, the contact and OPE diagrams contribute to the scattering amplitudes. At NLO, there are one-loop corrections to these LO diagrams and also TPE diagrams. The calculated scattering amplitudes for these diagrams are listed in Appendix A.1. The $D^{(*)}\bar{B}^{(*)}[0(0^{+})]$ coupled-channel potential in the coordinate space is
$\begin{pmatrix}V_{D\bar{B}-D\bar{B}}&V_{D\bar{B}-D^{*}\bar{B}^{*}}\\ V_{D^{*}\bar{B}^{*}-D\bar{B}}&V_{D^{*}\bar{B}^{*}-D^{*}\bar{B}^{*}}\end{pmatrix},$
where the inelastic scattering $D\bar{B}\to D^{*}\bar{B}^{*}$ and $D^{*}\bar{B}^{*}\to D\bar{B}$ appear as non-diagonal elements in the $2\times 2$ potential matrix. Then we solve the coupled-channel Schrödinger equation and obtain the numerical results including the binding energy $E$, and the probability of the individual channel $P_{i}$, which are shown in Table 10.
Here we consider the variation on the cutoff $\Lambda$ just as in the single-channel case. Compared with the single-channel $D\bar{B}[0(0^{+}]$ results in Table 9, we find coupled channels somewhat strengthen the binding energy: from 0.8 GeV in the single channel to 1.1 GeV in the coupled channels at $\Lambda=0.7$ GeV, for example. Also, because of the large mass gap (about $190~{}{\rm{MeV}}$) between the $D\bar{B}$ and $D^{*}\bar{B}^{*}$ thresholds, the contribution of $D^{*}\bar{B}^{*}(^{1}S_{0})$ is $0.1\%$ at $\Lambda=0.5~{}\rm{GeV}$. If we raise the cutoff to $0.7~{}\rm{GeV}$ and $1.0~{}\rm{GeV}$, the probability of $D^{*}\bar{B}^{*}(^{1}S_{0})$ increases to $0.3\%$ and $0.9\%$, respectively.
For the $D^{(*)}\bar{B}^{(*)}[0(1^{+})]$ state, all the coupled-channel diagrams and corresponding scattering amplitudes are shown in Appendix A.1-A.5. Using these amplitudes, we can obtain a $3\times 3$ coupled-channel potential matrix,
$\begin{pmatrix}V_{D\bar{B}^{*}-D\bar{B}^{*}}&V_{D\bar{B}^{*}-D^{*}\bar{B}}&V_{D\bar{B}^{*}-D^{*}\bar{B}^{*}}\\ V_{D^{*}\bar{B}-D\bar{B}^{*}}&V_{D^{*}\bar{B}-D^{*}\bar{B}}&V_{D^{*}\bar{B}-D^{*}\bar{B}^{*}}\\ V_{D^{*}\bar{B}^{*}-D\bar{B}^{*}}&V_{D^{*}\bar{B}^{*}-D^{*}\bar{B}}&V_{D^{*}\bar{B}^{*}-D^{*}\bar{B}^{*}}\end{pmatrix}.$
We still take the value of the cutoff parameter at $0.5~{}\rm{GeV}$, $0.7~{}\rm{GeV}$, and $1.0~{}\rm{GeV}$ to evaluate binding energy and probability $P_{i}$, respectively. The numerical results are listed in Table 11. We find no bound state in the $D^{(*)}\bar{B}^{(*)}[0(1^{+})]$ system which is consistent with the single-channel $D\bar{B}^{*}[0(1^{+})]$ calculation. The dominant channel is $D\bar{B}^{*}(^{3}S_{1})$, with a probability of $99.9\%-99.5\%$.
According to the above analyses, we can find that the coupled-channel effects do not change our main conclusions obtained in the single-channel case.
Table 10: The numerical results for $D^{(*)}\bar{B}^{(*)}[0(0^{+})]$ systems in $S$-wave. The cutoff parameter $\Lambda$, and binding energy $E$ relative to the threshold of $D\bar{B}$ are in units of $\rm{GeV}$, and $\rm{MeV}$ respectively. $\cdots$” means that there is no bound solution.
      $\Lambda$        $0.5$        $0.7$        $1.0$
      $E$       $\cdots$       $1.1$       $7.3$
      $P_{1}$       $99.9\%$       $99.7\%$       $99.1\%$
      $P_{2}$       $0.1\%$       $0.3\%$       $0.9\%$
Table 11: The numerical results for $D^{(*)}\bar{B}^{(*)}[0(1^{+})]$ systems in $S$-wave. The cutoff parameter $\Lambda$, and binding energy $E$ relative to the threshold of $D\bar{B}^{*}$ are in units of $\rm{GeV}$, and $\rm{MeV}$ respectively. $\cdots$” means that there is no bound solution.
      $\Lambda$        $0.5$        $0.7$        $1.0$
      $E$       $\cdots$       $\cdots$       $\cdots$
      $P_{1}$       $99.9\%$       $99.7\%$       $99.5\%$
      $P_{2}$       $0.1\%$       $0.3\%$       $0.5\%$
      $P_{3}$       $0.0\%$       $0.0\%$       $0.0\%$
IV.3 Two-body scattering of $D^{(*)}\bar{B}^{(*)}$ systems
In addition to the binding energies, we further derive information about $D^{(*)}\bar{B}^{(*)}$ elastic scattering by solving the Lippmann-Schwinger equation. In quantum theory, the scattering rate represents the probability per unit time for a scattering event, it is proportional to the cross section, the number of targets, and the flux Thomson:2013zua . Using Eqs. (12)-(16), we now calculate the scattering rates between $D^{(*)}$ and $\bar{B}^{(*)}$. The variations of the scattering rates $k\sigma(k)$ with cutoffs are shown in Figs. 9 and 10, as functions of the center-of-mass energy.